# Towards Robust Dialogue Breakdown Detection: Addressing Disruptors in Large Language Models with Self-Guided Reasoning

### v1

@@ Abstract
Large language models (LLMs) are rapidly changing various domains. However, their capabilities in handling conversational breakdowns still require an in-depth exploration. This paper addresses the challenge of detecting and mitigating dialogue breakdowns within LLM-driven conversational systems. While powerful models from OpenAI and Anthropic excel in many dialogue tasks, they can still produce incoherent or contradictory responses, commonly referred to as breakdowns, which undermine user trust. To tackle this, we propose an approach that combines specialized fine-tuning with advanced prompting strategies, including few-shot learning, chain-of-thought reasoning, and analogical prompting. In particular, we fine-tune a small 8B model and demonstrate its robust classification and calibration capabilities in English and Japanese dialogue. We also validate its generalization on the BETOLD dataset, achieving a 7% accuracy improvement over its base model. Furthermore, we introduce a real-time deployment architecture that selectively escalates suspicious responses to more resource-intensive frontier models only when breakdowns are detected, significantly cutting operational expenses and energy consumption. Experimental results show our method surpasses prior state-of-the-art specialized classifiers while also narrowing performance gaps between smaller open-source models and large proprietary ones. Our approach offers a scalable solution for robust conversational AI in high-impact domains by combining efficiency, interpretability, and reliability.
{IEEEImpStatement}
Our research advances conversational AI by reducing incoherent or contradictory responses, known as dialogue breakdowns, which erode user confidence. Unlike previous approaches dependent on large, specialized models, we achieve state-of-the-art accuracy and strong calibration using a fine-tuned 8B-parameter language model. This improvement significantly cuts operational costs while maintaining reliability in detecting breakdowns in multilingual and task-oriented dialogues. We further reduce compute demands by calling on larger, more resource-intensive models only when absolutely necessary. This balance between scalability, accuracy, and sustainability facilitates the greater adoption of safe, trustworthy AI systems in critical areas such as customer support, education, and healthcare. By aligning system confidence with human judgment, our framework promotes greater transparency and user acceptance, paving the way for more robust and accessible deployment of next-generation conversational technologies.
{IEEEkeywords}
large language models, conversational artificial intelligence, human-computer interaction
1 Introduction
Conversational artificial intelligence (AI) has experienced exponential growth driven by the rapid advancement and widespread adoption of large language models (LLMs) such as ChatGPT [1], Claude [2], and Llama [3]. The unprecedented success of ChatGPT, which attracted one million users in five days and exceeded 100 million in two months [4, 5], emphasizes both the unique capabilities and the immense potential of LLMs. These generative models have rapidly permeated numerous sectors, including customer service, content creation, marketing, and education, fundamentally redefining human-AI interactions [6, 7, 8].
Table 1: Sample Dialogue
@@ 1 Introduction
   Determine if the following utterance leads to a dialogue breakdown:
@@
Despite these advancements, the rapid integration of LLMs into critical domains has surfaced significant challenges, notably dialogue breakdowns, which can severely disrupt user trust and conversational effectiveness [9, 10, 11]. Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [12, 13]. Table 1 illustrates an example scenario where the assistant’s initial and subsequent utterances contradict each other, illustrating a common breakdown scenario.
The importance of addressing dialogue breakdowns grows alongside adopting LLMs in high-stakes environments, such as customer support and educational applications [6, 7, 8]. Moreover, the tendency of LLMs to produce overly confident yet potentially inaccurate or hallucinated responses further complicates their reliable deployment [14, 15]. Consequently, robust methodologies are clearly needed to detect and mitigate dialogue breakdowns to preserve conversational quality and user trust.
@@
Addressing these challenges, we introduce innovative approaches that take advantage of generalist LLMs’ reasoning capabilities through supervised fine-tuning and structured prompting. Specifically, we fine-tune the efficient and powerful Llama-3.1 8B model [3] on both English and Japanese tracks from the DBDC5 dataset, incorporating synthetic reasoning trajectories generated by the superior Llama-3.3 70B model. These distilled reasoning tracks guide and enhance interpretability during training, significantly improving the model’s decision-making process. We rigorously evaluate our model’s generalization capability using the challenging BETOLD dataset [19], which explicitly focuses on task-oriented dialogue breakdowns across diverse conversational contexts.
Additionally, we provide a comprehensive comparative analysis of both closed-source frontier models (OpenAI [1] and Anthropic [2]) and open-source alternatives (Meta [3], Mistral AI [20], and DeepSeek [21]), specifically targeting their dialogue breakdown detection capabilities. We explore advanced prompting methodologies including few-shot learning [22, 23], zero-shot chain-of-thought prompting [24], and propose analogical reasoning enhanced by curricular learning strategies [25, 26, 27]. These strategies encourage systematic reasoning and adaptability without extensive labeled datasets.
Through these novel techniques, we establish state-of-the-art benchmarks for the Dialogue Breakdown Detection Challenge. Furthermore, we examine calibration metrics alongside traditional classification metrics, gaining insights into model overconfidence and reliability. Our proposed architecture introduces a cost-effective, real-time dialogue breakdown monitoring system that utilizes a fine-tuned Llama-3.1 8B model, selectively invoking resource-intensive models such as GPT-4, DeepSeek-R1, Sonnet-3.5, or Llama-3.1 405B when necessary. Given that high-capacity models significantly raise operational costs and energy consumption, our architecture optimizes resource utilization, achieving substantial cost reductions and promoting sustainability.
@@
  • Conducted the first systematic comparative analysis between open-source and frontier closed-source models, setting new benchmarks in dialogue breakdown detection.
@@ 2 Related Work
2.1 Specialized Models for Dialogue Breakdown Detection
Research into dialogue breakdown detection has yielded specialized classifiers that push the performance on datasets like DBDC5 (see Section 5.1 for more information) [12]. The top-performing approaches leverage pre-trained transformer [30] encoders fine-tuned for breakdown classification. For example, the best model on the DBDC5 English track was an augmented BERT-based classifier, BERT+SSMBA [31, 32]. BERT+SSMBA incorporates unlabeled dialogue data via two complementary techniques: extended pre-training on dialogue-rich datasets such as Reddit and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [33]. This approach exploits unlabeled data to improve the classifier’s robustness in detecting dialogue breakdowns.
@@ 2.1 Specialized Models for Dialogue Breakdown Detection
2.2 Current State of Conversational Agents
Beyond detecting breakdowns, state-of-the-art conversational agents incorporate strategies to mitigate breakdowns. For example, if a breakdown is detected, a system might employ a recovery strategy (as studied in DBDC5’s recovery track) to ask for clarification or provide corrections [12]. Leading chatbots like ChatGPT and Claude are also trained via alignment techniques such as reinforcement learning from human feedback (RLHF) to minimize toxic, incoherent, or nonsensical outputs that could derail a dialogue [1, 2]. Models like Claude-2 even engage in internal debates during training to identify and eliminate reasoning flaws, aiming to reduce the chance of breakdowns due to factual errors or contradictions [37]. The net result is that modern conversational AI systems are progressively improving at maintaining coherent dialogues despite ambiguous user inputs.
2.3 Techniques in Conversational AI
@@
Analogical Reasoning. This technique guides a language model to draw on relevant past examples or scenarios by analogy when confronting a new problem. Instead of relying solely on provided examples, the model generates analogous examples as context. Inspired by human problem-solving via analogy, analogical prompting has been shown to improve reasoning accuracy [25, 38]. It ”prompts language models to self-generate relevant exemplars or knowledge in context, before proceeding to solve the given problem,” removing the need for hand-crafted exemplars and tailoring the reasoning to the specific query [25]. In practice, a dialogue agent using analogical reasoning might recall a similar conversation or scenario it has seen and use that to form a better response to a user query.
@@
Chain-of-Thought Reasoning. Chain-of-thought (CoT) is a reasoning technique in which the model is encouraged to produce explicit intermediate reasoning steps (”thinking aloud”) before giving a final answer. This has effectively improved logical consistency and arithmetic or commonsense reasoning in LLMs [39, 40, 24]. For dialogue agents, chain-of-thought can be used internally (the model can reason through the user’s query, context, and knowledge base step-by-step) to formulate a correct and context-appropriate response. This reduces mistakes and non-sequiturs. CoT reasoning is now a common technique to bolster LLM robustness on tasks requiring multi-step inference or clarification, as it significantly ”enhances their ability to reason” about the conversation [24, 40, 39].
@@
Zero-Shot and Few-Shot Learning. Unlike traditional dialogue systems requiring extensive task-specific training, LLM-based agents excel at in-context learning as they can adapt to new instructions or domains given few examples or none at all. In few-shot learning, a handful of example dialogues or question-answer pairs are provided in the prompt, and the model generalizes the pattern to the new query. The seminal GPT-3 work showed that large models can perform new language tasks by observing just a few demonstrations, essentially treating ”language models as few-shot learners.” [22]. This allows a conversational agent to be quickly customized to a new domain by giving it a few sample conversations as guidance rather than retraining the model. LLMs can often do this for dialogue tasks. For instance, a well-designed prompt can instruct the model to play the role of a customer service agent and will generate reasonable dialogue without any fine-tuning. In-context learning empowers LLM-based agents to handle various conversational scenarios with minimal additional training.
3 Problem Definition
@@
Dialogue breakdown, defined as the deterioration of coherence, relevance, or conversational fluidity between a user and a conversational agent, remains a significant challenge in conversational AI[12, 29]. Such breakdowns may manifest as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that hinder the natural progression of dialogue and consequently degrade user trust and satisfaction [10, 11]. As LLMs from the OpenAI [1], Claude [2], and Llama [3] families are rapidly adopted for diverse conversational tasks, addressing these dialogue disruptions has become paramount. Moreover, dialogue breakdown detection becomes even more critical due to LLMs’ propensity for confident but potentially incorrect or hallucinated outputs, resulting in compounded user frustration [14, 15].
@@
$D=\bigl{(}u_{1},s_{1},\;u_{2},s_{2},\;\dots,\;u_{n},s_{n}\bigr{)}$
@@
3.1 Utterance-Level Breakdown Detection
@@
$\mathcal{H}_{i}=\bigl{(}u_{1},s_{1},\dots,u_{i-1},s_{i-1},u_{i}\bigr{)}$
@@
3.2 Consolidation of Three-Class Annotations
Previous datasets like DBDC [12] use three-level labels: Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB). Given the subjectivity of human annotations in dialogue breakdown detection, we often consolidate PB into B. When multiple human annotators supply these labels, let $p_{i}$ be the fraction who labeled $s_{i}$ as B or PB. We can convert each utterance into a binary label $b_{i}$ by thresholding:
@@ 3.2 Consolidation of Three-Class Annotations
here $p({b_{i}}\mid s_{i})$ is the fraction of human annotators labeling $s_{i}$ as a breakdown. Throughout this paper, “breakdown vs. non-breakdown” refers to this binary mapping.
3.3 Conversation-Level Labeling
@@ 3.4 Confidence Calibration
4 Methodology
Our approach combines specialized fine-tuning and advanced prompting to achieve robust dialogue breakdown detection in both open-domain and task-oriented conversations. We focus on three key components: a compact fine-tuned model to recognize dialogue breakdowns efficiently, leveraging advanced prompting strategies for generalist LLMs, and propose a multi-tier inference architecture that calls upon frontier models only when necessary for corrections, optimizing both cost and accuracy.
4.1 Supervised Fine-tuning with Reasoning Augmentation
To deploy an efficient alternative for real-time monitoring, we fine-tune a smaller model on labeled breakdown data using supervised fine-tuning (SFT). We choose Llama-3.1 8B [3] as a balance between accuracy and computational cost. The model is fine-tuned using the DBDC5 English and Japanese tracks [12] for per-utterance breakdown labels. Let:
@@ 4.1 Supervised Fine-tuning with Reasoning Augmentation
We augment the original training input with this synthetic reasoning, effectively creating a more informative training set $\mathcal{D}^{\prime}$. We fine-tune our student model $S$ to minimize the cross-entropy loss ($\mathcal{L}_{CE}$) [41] between predicted and true labels:
@@
here, $y_{i}$ denotes the true label for utterance $i$, and $\hat{y}_{i}$ denotes the predicted probability of breakdown. The phase allows the student to see both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s structured reasoning patterns $r_{i}$ during training, learning to distill these insights.
4.2 Advanced Prompting Strategies

### v2

# Detect, Explain, Escalate: Low-Carbon Dialogue Breakdown Management for LLM-Powered Agents

@@ Abstract
While Large Language Models (LLMs) are transforming numerous applications, their susceptibility to conversational breakdowns remains a critical challenge undermining user trust. This paper introduces a “Detect, Explain, Escalate” framework to manage dialogue breakdowns in LLM-powered agents, emphasizing low-carbon operation. Our approach integrates two key strategies: (1) We fine-tune a compact 8B-parameter model, augmented with teacher-generated reasoning traces, which serves as an efficient real-time breakdown ‘detector’ and ‘explainer.’ This model demonstrates robust classification and calibration on English and Japanese dialogues, and generalizes well to the BETOLD dataset, improving accuracy by 7% over its baseline. (2) We systematically evaluate frontier LLMs using advanced prompting (few-shot, chain-of-thought, analogical reasoning) for high-fidelity breakdown assessment. These are integrated into an ‘escalation’ architecture where our efficient detector defers to larger models only when necessary, substantially reducing operational costs and energy consumption. Our fine-tuned model and prompting strategies establish new state-of-the-art results on dialogue breakdown detection benchmarks, outperforming specialized classifiers and significantly narrowing the performance gap to larger proprietary models. The proposed monitor-escalate pipeline reduces inference costs by 54%, offering a scalable, efficient, and more interpretable solution for robust conversational AI in high-impact domains. Code and models will be publicly released.
Index Terms:
large language models, conversational artificial intelligence, human-computer interaction, dialogue systems, model calibration
I Introduction
Conversational artificial intelligence (AI) has experienced exponential growth driven by the rapid advancement and widespread adoption of large language models (LLMs) such as ChatGPT [1], Claude [2], and Llama [3]. The unprecedented success of ChatGPT, which attracted one million users in five days and exceeded 100 million in two months [4, 5], highlights both the unique capabilities and the immense potential of LLMs. These generative models have rapidly permeated numerous sectors, including customer service, content creation, marketing, and education, fundamentally reshaping human-AI interactions [6, 7, 8].
TABLE I: Sample dialogue illustrating contradictory turns.
@@ I Introduction
Determine if the next utterance causes a breakdown:
@@
Despite these advancements, the rapid integration of LLMs into critical domains has surfaced significant challenges, notably dialogue breakdowns, which can severely disrupt user trust and conversational effectiveness [9, 10, 11]. Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [12, 13]. Table I presents an example where the assistant’s utterances are contradictory, illustrating a common breakdown scenario.
Addressing dialogue breakdowns becomes increasingly critical with the increasing adoption of LLMs in high-stakes environments [6, 7, 8]. Moreover, the tendency of LLMs to produce overly confident yet potentially inaccurate or hallucinated responses further complicates their reliable deployment [14, 15]. Consequently, there is a pressing need for robust methodologies to detect and mitigate dialogue breakdowns, thereby preserving conversational quality and user trust.
@@
Addressing these challenges, we introduce approaches that leverage the reasoning capabilities of generalist LLMs through supervised fine-tuning and structured prompting. Specifically, we fine-tune the efficient and powerful Llama-3.1 8B model [3] on both English and Japanese tracks from the DBDC5 dataset, incorporating synthetic reasoning trajectories generated by a larger, more capable Llama-3.3 70B model. These distilled reasoning tracks, incorporated during fine-tuning, aim to improve the student model’s decision-making process and enhance the interpretability of its predictions. We rigorously evaluate its generalization capability using the challenging BETOLD dataset [19], which explicitly focuses on task-oriented dialogue breakdowns across diverse conversational contexts.
Additionally, we provide a comprehensive comparative analysis of both closed-source frontier models (OpenAI [1] and Anthropic [2]) and open-source alternatives (Meta [3], Mistral AI [20], and DeepSeek [21]), specifically targeting their dialogue breakdown detection capabilities. We explore advanced prompting methodologies including few-shot learning [22, 23], zero-shot chain-of-thought prompting [24], and propose analogical reasoning enhanced by curricular learning strategies [25, 26, 27]. These strategies encourage systematic reasoning and adaptability, often reducing the need for extensive labeled datasets.
Through these novel techniques, we establish state-of-the-art benchmarks for the Dialogue Breakdown Detection Challenge. Furthermore, we examine calibration metrics alongside traditional classification metrics, gaining insights into model overconfidence and reliability. Our proposed architecture introduces a cost-effective, real-time dialogue breakdown monitoring system that utilizes a fine-tuned Llama-3.1 8B model, selectively invoking resource-intensive models such as GPT-4, DeepSeek-R1, Sonnet-3.5, or Llama-3.1 405B when necessary. Our architecture optimizes resource utilization by strategically employing high-capacity models, which typically incur significant operational costs and energy consumption, only when necessary. This approach achieves substantial cost reductions and promotes sustainability.
@@
  • Conducted the first comprehensive comparative analysis across a wide range of open-source and frontier closed-source models, establishing new benchmarks in dialogue breakdown detection.
@@ II Related Work
II-A Specialized Models for Dialogue Breakdown Detection
Research into dialogue breakdown detection has yielded specialized classifiers that push the performance on datasets like DBDC5 (see Section V-A for more information) [12]. The top-performing approaches leverage pre-trained transformer [30] encoders fine-tuned for breakdown classification. For example, the best model on the DBDC5 English track was an augmented BERT-based classifier, BERT+SSMBA [31, 32]. BERT+SSMBA incorporates unlabeled dialogue data via two complementary techniques: extended pre-training on dialogue-rich datasets such as Reddit and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [33]. This approach exploits unlabeled data to improve the classifier’s robustness in detecting dialogue breakdowns.
@@ II-A Specialized Models for Dialogue Breakdown Detection
II-B Current State of Conversational Agents
Beyond detecting breakdowns, state-of-the-art conversational agents incorporate strategies to mitigate breakdowns. For example, if a breakdown is detected, a system might employ a recovery strategy (as studied in DBDC5’s recovery track) to ask for clarification or provide corrections [12]. Leading chatbots like ChatGPT and Claude are also trained via alignment techniques such as reinforcement learning from human feedback (RLHF) to minimize toxic, incoherent, or nonsensical outputs that could derail a dialogue [1, 2]. Models like Claude-2 even engage in internal debates during training to identify and eliminate reasoning flaws, aiming to reduce the likelihood of breakdowns due to factual errors or contradictions [37]. The net result is that modern conversational AI systems are progressively improving at maintaining coherent dialogues despite ambiguous user inputs.
II-C Techniques in Conversational AI
@@
Analogical Reasoning. This technique guides a language model to draw on relevant past examples or scenarios by analogy when confronting a new problem. Instead of relying solely on provided examples, the model generates analogous examples as context. Inspired by human problem-solving via analogy, analogical prompting has been shown to improve reasoning accuracy [25, 38]. It “prompts language models to self-generate relevant exemplars or knowledge in context, before proceeding to solve the given problem,” thereby removing the need for hand-crafted exemplars and tailoring the reasoning to the specific query [25]. In practice, a dialogue agent using analogical reasoning might recall a similar conversation or scenario it has encountered and use that information to form a better response to a user query.
@@
Chain-of-Thought Reasoning. Chain-of-thought (CoT) is a reasoning technique in which the model is encouraged to produce explicit intermediate reasoning steps (“thinking aloud”) before giving a final answer. This has effectively improved logical consistency and arithmetic or commonsense reasoning in LLMs [39, 40, 24]. For dialogue agents, chain-of-thought can be used internally (the model can reason through the user’s query, context, and knowledge base step-by-step) to formulate a correct and context-appropriate response. This reduces mistakes and non-sequiturs. CoT reasoning is now a common technique to bolster LLM robustness on tasks requiring multi-step inference or clarification, as it significantly “enhances their ability to reason” about the conversation [24, 40, 39].
@@
Zero-Shot and Few-Shot Learning. Unlike traditional dialogue systems requiring extensive task-specific training, LLM-based agents excel at in-context learning as they can adapt to new instructions or domains given few examples or none at all. In few-shot learning, a handful of example dialogues or question-answer pairs are provided in the prompt, and the model generalizes the pattern to the new query. The seminal GPT-3 work showed that large models can perform new language tasks by observing just a few demonstrations, essentially treating “language models as few-shot learners” [22]. This allows a conversational agent to be quickly customized to a new domain by giving it a few sample conversations as guidance rather than retraining the model. LLMs can often do this for dialogue tasks. For instance, a well-designed prompt can instruct the model to play the role of a customer service agent and will generate reasonable dialogue without any fine-tuning. In-context learning empowers LLM-based agents to handle various conversational scenarios with minimal additional training.
III Problem Definition
@@
Dialogue breakdown, defined as the deterioration of coherence, relevance, or conversational fluidity between a user and a conversational agent, remains a significant challenge in conversational AI[12, 29]. Such breakdowns may manifest as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that hinder the natural progression of dialogue and consequently degrade user trust and satisfaction [10, 11]. As LLMs from the OpenAI [1], Claude [2], and Llama [3] families are rapidly adopted for diverse conversational tasks, addressing these dialogue disruptions has become paramount. Moreover, dialogue breakdown detection becomes even more critical due to LLMs’ propensity for confident but potentially incorrect or hallucinated outputs, resulting in increased user frustration [14, 15].
@@
$D=\bigl{(}u_{1},s_{1},\;u_{2},s_{2},\;\dots,\;u_{n},s_{n}\bigr{)}.$
@@
III-A Utterance-Level Breakdown Detection
@@
$\mathcal{H}_{i}=\bigl{(}u_{1},s_{1},\dots,u_{i-1},s_{i-1},u_{i}\bigr{)}.$
@@
III-B Consolidation of Three-Class Annotations
Previous datasets like DBDC [12] use three-level labels: Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB). Given the subjectivity of human annotations in dialogue breakdown detection, we consolidate ‘Possible Breakdown’ PB into the ‘Breakdown’ (B) class. When multiple human annotators supply these labels, let $p_{i}$ be the fraction who labeled $s_{i}$ as B or PB. We can convert each utterance into a binary label $b_{i}$ by thresholding:
@@ III-B Consolidation of Three-Class Annotations
where $p({b_{i}}\mid s_{i})$ is the fraction of human annotators labeling $s_{i}$ as a breakdown. Throughout this paper, the binary classification of “breakdown versus non-breakdown” refers to this mapping.
III-C Conversation-Level Labeling
@@ III-D Confidence Calibration
IV Methodology
Our approach combines specialized fine-tuning and advanced prompting to achieve robust dialogue breakdown detection in both open-domain and task-oriented conversations. We focus on three key components: a compact fine-tuned model to detect dialogue breakdowns efficiently, leveraging advanced prompting strategies for generalist LLMs, and a proposed multi-tier inference architecture that calls upon frontier models only when necessary for corrections, optimizing both cost and accuracy.
IV-A Supervised Fine-tuning with Reasoning Augmentation
To deploy an efficient alternative for real-time monitoring, we fine-tune a smaller model on labeled breakdown data using supervised fine-tuning (SFT). We choose Llama-3.1 8B [3] as it offers a good balance between accuracy and computational cost. The model is fine-tuned using the DBDC5 English and Japanese tracks [12] for per-utterance breakdown labels. Let:
@@ IV-A Supervised Fine-tuning with Reasoning Augmentation
We augment the original training input with this synthetic reasoning, effectively creating a more informative training set $\mathcal{D}^{\prime}$. We fine-tune our student model $S$ on the augmented dataset $\mathcal{D}^{\prime}$. The model is trained to predict the binary label $b_{i}$ (minimizing cross-entropy loss $\mathcal{L}_{CE}$) and to generate a textual justification $\hat{j}_{i}$. For the latter, the teacher-generated reasoning traces $r_{i}$ serve as training targets, enabling the student model to distill these structured reasoning patterns for improved classification and explanation. The cross-entropy loss [41] is defined as:
@@
where, $y_{i}$ denotes the true label for utterance $i$, and $\hat{y}_{i}$ denotes the predicted probability of breakdown. This training phase allows the student model to observe both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s structured reasoning patterns $r_{i}$ during training, learning to distill these insights.
IV-B Advanced Prompting Strategies